### v1

@@ IV-B Sampling Skew Diagnosis
In Line 11 of Alg. 2, the differences in $x$ and $y$ coordinates between consecutive robot poses are used to calculate trajectory orientations, notated as $\rho_{i}=\arctan(\Delta y_{i}/\Delta x_{i})$ and $\rho_{i+1}=\arctan(\Delta y_{i+1}/\Delta x_{i+1})$. Then variation of trajectory orientations $\Delta\rho_{i}=\rho_{i+1}-\rho_{i}$ is then normalized to the range $[-\pi,\pi]$ (in Line 12 of Alg. 2). If the absolute value $|\Delta\rho_{i}|$ exceeds the threshold $\eta$, the corresponding pose $p_{i}$ is identified as a high-resistance point and added to the set $\mathcal{H}$. Finally, the algorithm returns the set $\mathcal{H}$, representing the high-resistance area. The key intuition behind Alg. 2 is to capture regions where the meta planner shows high uncertainty in challenging scenarios, indicated by a high variance in the trajectory.
Note that there are two technical aspects that merit particular attention: ❶ Alg. 2 employs variations in trajectory orientations $\Delta\rho$ rather than robot orientations $\Delta w$ as the metric for comparison against the threshold $\eta$. This design effectively captures behaviors such as back-and-forth movements, where the robot’s orientation may remain relatively unchanged and thus cannot be detected by $\Delta w$. However, these movements are also clear indicators of challenges faced by the planner. ❷ Alg. 2 strategically excludes failure trajectories. This exclusion ensures that any points of high resistance can ultimately reach the goal. Allowing the robot to persist in areas that are excessively challenging would cause the planner to become indefinitely stuck, thereby introducing noise into the training of the meta-planner. This process allows for the identification of challenging areas where the robot experiences significant difficulties, enabling targeted interventions to improve the planner’s performance in these regions.
@@
IV-C Sampling Skew Mitigation
High-Resistance Area Up-sampling. In Lines 10-11 of Alg. 1, the process of high-resistance area up-sampling is implemented to enhance the robustness and adaptability of the meta-planner during training. This algorithm leverages threshold $\lambda\in[0,1]$ to determine whether to sample the robot’s initial position from the set of high-resistance areas $\mathcal{H}$. It ensures that the meta-planner frequently encounters and learns to navigate through high-resistance regions, thereby improving its ability to handle challenging scenarios. By balancing this targeted exploration with a general exploration of the environment, the meta-planner will train more efficiently and effectively in high-resistance areas. Additionally, $\mathcal{H}$ is updated iteratively in Line 20 of Alg. 1, ensuring that the set of high-resistance areas remains relevant and accurately reflects the current sampling challenges. This iterative update is crucial for maintaining the effectiveness of the up-sampling strategy. Overall, the up-sampling of high-resistance areas is critical for the meta-planner’s training as it enhances the planner’s robustness, promotes better generalization to new environments, and addresses sampling skew.
@@ V Evaluation
Table I: Navigation performance of Digimon and baseline methods under Same-Env setup.
Metric $\Rightarrow$ Method $\Downarrow$ Easy Medium Difficult
@@
Table II: Navigation performance of Digimon and baselines under Cross-Env and Cross-Level setups. The column names represent the test environment levels: All (across all three levels), Easy, Medium and Difficult.
Training Env. Metric $\Rightarrow$ Method $\Downarrow$ All Easy Medium Difficult
@@ V-A Experiment Setup
Simulation & Planners. Following existing works [xiao2020appld, xu2021applr], we implement Digimon using the widely-used Gazebo [Gazebo] simulator with a ClearPath Jackal differential drive ground robot as a robot model. Due to space limit, we leave the details of the classical planner and the meta-planner that Digimon builds upon for evaluation to Appendix VII.
Baselines & Dataset. We compare Digimon with two widely used classical planners (DWA [fox1997dwa] and DWA-Fast [dwafast]), one representative static meta-planner, BayesOpt [snoek2012practical] and one SOTA dynamic RL-based meta-planner APPLR [xu2021applr]. We choose the same dataset as APPLR, namely BARN [perille2020benchmarking] dataset, as it includes various maps for robot navigation inside the Gazebo simulator and it has different difficulty levels’ map (i.e., Easy, Medium, and Difficult) to evaluate. This allows for a convenient evaluation of various planners’ performance across these varied settings.
@@
Comparison Setup. We introduce 3 evaluation setups: (1) Same-Env: The meta-planner is trained and tested in the same environment set with predefined but different start positions to avoid overfitting. (2) Cross-Env: The meta-planner is trained and tested at the same difficulty level but in different environments. (3) Cross-Level: The meta-planner is trained on one difficulty level but tested on a different difficulty level.
V-B Same-environment Performance
We evaluate Digimon and four baselines under the Same-Env setup. DWA, DWA-Fast, and BayesOpt are static planners, so we directly implement their fixed configurations from the training environment and test them on the same environments with random starting positions. For APPLR and Digimon, we train the meta-planner policy on the training set and deploy it on the test environments.
@@ V-B Same-environment Performance
Table III: Performance of Digimon on PPO-based meta-planner.
PPO + APPLR PPO + Ours
NS $\uparrow$ 17.58 20.61
ATT $\downarrow$ 24.69 22.07
SR $\uparrow$ 55.56 65.28
CR $\downarrow$ 26.39 15.28
TR $\downarrow$ 18.06 19.44
Table IV: Performance of Digimon w/ and wo/ failure filtering.
@@
Table I shows that Digimon consistently achieves the highest NS score across all difficulty levels, demonstrating its ability to identify and improve performance in underrepresented scenarios. While DWA performs better on SR and CR in easy and medium environments, Digimon excels in the comprehensive NS metric. DWA’s conservative design prioritizes safety over time efficiency, leading to longer navigation times, as seen in reduced ATT and lower NS scores, making it less practical for navigation tasks. Since NS is a widely-used metric [xu2021applr, xiao2020appld], Digimon’s higher NS score effectively demonstrates its superiority over both static planners and the SOTA dynamic meta-planner.
V-C Cross-environment & Cross-level Performance
In this section, we compare Digimon and selected baselines under the Cross-Env and Cross-Level setup, as shown in the diagonal sub-tables in Table II. Under the Cross-Env setup, Digimon consistently achieves the highest NS score across all three environments, indicating its effectiveness in identifying and addressing underrepresented scenarios to enhance planner performance. For the Cross-Level experiments, Digimon excels in generalization by achieving the highest NS scores when the trained meta-planner is tested across different environmental difficulty levels. This demonstrates Digimon’s robustness and capability to maintain navigation efficacy in new and varying conditions. Additionally, when meta-planners trained in difficult environments are applied to easier ones, Digimon outperforms even those baselines specifically trained for simpler environments, while this trend is not observed in the baseline methods. This result suggests that the meta-planners trained by our method can learn more sophisticated and adaptable configurations, enabling Digimon to excel in environments that are less challenging.
V-D Diagnosis Effectiveness
To better understand why our diagnosis and mitigation strategy is effective, we introduce a variant called RandomSampling (RS). Unlike Digimon, which identifies high-resistance areas with Alg. 2, RS indiscriminately samples from all locations reached by the robot during training without strategic guidance, while all other designs remain consistent with Digimon. Results in Fig. 5 show that Digimon significantly outperforms RS, with higher NS scores in all environments. The reasons are two-fold. First, RS fails to expose the robot to challenging, hard-to-reach scenarios, limiting effective training in these high-resistance areas, which is a crucial aspect for satisfactory navigation performance. Second, as shown in Fig. 2, medium and hard environments have more complicated maps with many hard-to-reach scenarios and obstacles than the easy level. Random sampling in such environments often leads to the robot encountering “dead-corner” scenarios – areas that are hard to exit, leading to extended, unrewarding training trajectories. Such trajectories will destabilize the training of the RL policy in the meta-planner and degrade navigation performance. This analysis verifies the necessity of our high-resistance area up-sampling step in Digimon. By strategically selecting those hard-to-reach areas, we enhance the robot’s ability to perform well across a variety of navigational scenarios. Appendix VIII shows more results in the comparison between RS and Digimon.
(a) Easy
(b) Medium
(c) Difficult
Figure 5: Comparison of RandomSampling and Digimon. The meta-planners are trained on three different difficulty levels and tested across all levels.
V-E Efficiency
In this section, we evaluate the performance of different meta-planner policy checkpoints during training under the Cross-Env setup when Digimon is applied in a medium-level environment. Fig. 4 shows the average NS metric when those checkpoints are evaluated in testing environments. Digimon consistently outperforms APPLR, achieving higher navigation scores at all training stages with the same training steps. Furthermore, APPLR requires over four times training steps to achieve comparable performance of Digimon, and our method ultimately outperforms APPLR by 13.5%. These results demonstrate that our method significantly enhances the training efficiency of the RL policy within the meta-planner and improves the overall navigation performance of the robot. Besides NS, we also show more metrics of different meta-planner checkpoints during training (in Fig. 9 in Appendix).
V-F Plug-and-Play Applicability on Other RL Algorithm
In Sections V-B and V-C, we have demonstrated that Digimon enhances navigation performance when paired with the off-policy algorithm TD3 [fujimoto2018addressing] based meta-planner in APPLR. To further validate the versatility of Digimon, we extended it to another SOTA on-policy RL algorithm, Proximal Policy Optimization (PPO) [ppo]. For a fair comparison, we replace the TD3 training algorithm in APPLR with PPO. Results are shown in Table V-B. We can observe that Digimon outperforms baseline methods when applied to PPO, where four out of five metrics get significantly improved, and a negligible drop in timeout rate. This indicates that compared to APPLR, our diagnostic method is robust across various training algorithms and not limited to a specific algorithm.
V-G Other Experiments
@@
Ablation study. We conduct the ablation study to justify the importance of key components in Digimon and the sensitivity test against the key hyper-parameters. Details can be found in Appendix IX-B.
@@ VI Conclusion
In this paper, we identify and analyze the challenges of integrating RL with classical planning in meta-planners. While this integration enhances adaptability and explainability, it introduces sampling skew due to the conservative nature of classical planners, limiting the RL meta-planners’ exploration and performance in complex environments.
To address these issues, we propose Digimon, a diagnostic and mitigation framework for RL-based meta-planners. Digimon analyzes robot behaviors, identifies exploration bottlenecks and applies a targeted up-sampling strategy to balance the sampling data, enabling more effective learning and exploration. Our experiments demonstrate that Digimon improves navigation performance by 13.5%+, enhances robustness in unseen environments, and quadruples training efficiency.
Limitations. Although Digimon greatly boosts efficiency, further optimization of training duration is possible. The current pipeline depends on executing the classical planner to collect data, meaning increased complexity of the classical planner can raise training costs. Additionally, Digimon is specifically designed for navigation tasks and may not apply to manipulation tasks, which we leave for future work.
@@ VII Experiment Setup
The meta-planner policy $\pi$ is trained to update the planning configurations $\theta$ of the classical planner. We maintain the same settings as APPLR, utilizing the DWA classical planner and the same action space $\theta$, including ${\tt max\_vel\_x}$, ${\tt max\_vel\_theta}$, ${\tt vx\_samples}$, ${\tt vtheta\_samples}$, ${\tt occdist\_scale}$, ${\tt pdist\_scale}$, ${\tt gdist\_scale}$, and ${\tt inflation\_radius}$ [^1] https://wiki.ros.org/dwa_local_planner We use the ROS dynamic reconfigure client to dynamically change planner parameters. The global goal $\alpha$ is assigned manually, while $\theta_{0}$ is the default set of parameters provided by the robot manufacturer. [^2] http://wiki.ros.org/costmap_2d Table V shows the details of each configuration of the meta-planner.
Configuration Min Max Meaning
${\tt max\_vel\_x}$ 0.1 2.0 m/s. Maximum threshold for linear velocity.
${\tt max\_vel\_theta}$ 0.314 3.14 rad/s. The absolute value of the maximum rotational velocity for the robot.
${\tt vx\_samples}$ 4 12 The number of samples to use when exploring the x velocity space.
${\tt vtheta\_samples}$ 8 40 The number of samples to use when exploring the theta velocity space
${\tt path\_distance\_bias}$ 0.1 0.5 The weight for how much the robot should stay close to the path it was given.
${\tt goal\_distance\_bias}$ 0.1 2 The weight for how much the robot should attempt to reach its local goal.
${\tt inflation\_radius}$ 0.1 0.6 Controls how far away the zero cost point is from obstacle.
Table V: Action Space of Meta-Planner Policy $\pi$.
VIII Diagnosis Effectiveness
In this section, we delve deeper into the comparative analysis between RandomSampling and Digimon by applying both methods across environments of varying complexity: Easy, Medium, and Difficult. We conducted separate experiments for each difficulty level, which allowed for detailed and direct comparisons. The results of these experiments are represented in Figures 67, and 8.
Consistent with the observations discussed in Section V-D, our Digimon continues to outperform RandomSampling in overall navigation performance. This superiority is quantified by a higher Navigation Score (NS), indicating Digimon’s effectiveness in handling complex navigation tasks. This comparison not only highlights the strengths of Digimon but also contributes to a better understanding of its operational dynamics across different environmental settings.

### v3

@@ IV-B Training Inefficiency Diagnosis
In Line 11 of Alg. 2, the differences in $x$ and $y$ coordinates between consecutive robot poses are used to calculate trajectory orientations, notated as $\rho_{i}=\arctan(\Delta y_{i}/\Delta x_{i})$ and $\rho_{i+1}=\arctan(\Delta y_{i+1}/\Delta x_{i+1})$. Then variation of trajectory orientations $\Delta\rho_{i}=\rho_{i+1}-\rho_{i}$ is then normalized to the range $\pi,\pi]$ (in Line 12 of Alg. 2). If the absolute value $|\Delta\rho_{i}|$ exceeds the threshold $\eta$, the corresponding pose $p_{i}$ is identified as a high-resistance point and added to the set $\mathcal{H}$. Finally, the algorithm returns the set $\mathcal{H}$, representing the high-resistance area. The key intuition behind Alg. 2 is to capture regions where the parameter tuner shows high uncertainty in challenging scenarios, indicated by a high variance in the trajectory.
Note that there are two technical aspects that merit particular attention: ❶ Alg. 2 employs variations in trajectory orientations $\Delta\rho$ rather than robot orientations $\Delta w$ as the metric for comparison against the threshold $\eta$. This design effectively captures behaviors such as back-and-forth movements, where the robot’s orientation may remain relatively unchanged and thus cannot be detected by $\Delta w$. However, these movements are also clear indicators of challenges faced by the planner. ❷ Alg. 2 strategically excludes failure trajectories. This exclusion ensures that any points of high resistance can ultimately reach the goal. Allowing the robot to persist in areas that are excessively challenging would cause the planner to become indefinitely stuck, thereby introducing noise into the training of the parameter tuner. This process allows for the identification of challenging areas where the robot experiences significant difficulties, enabling targeted interventions to improve the planner’s performance in these regions.
IV-C Training Inefficiency Mitigation
High-Resistance Area Up-sampling. In Lines 10-11 of Alg. 1, the process of high-resistance area up-sampling is implemented to enhance the robustness and adaptability of the parameter tuner during training. This algorithm leverages threshold $\lambda\in[0,1]$ to determine whether to sample the robot’s initial position from the set of high-resistance areas $\mathcal{H}$. It ensures that the parameter tuner frequently encounters and learns to navigate through high-resistance regions, thereby improving its ability to handle challenging scenarios. By balancing this targeted exploration with a general exploration of the environment, the parameter tuner will train more efficiently and effectively in high-resistance areas. Additionally, $\mathcal{H}$ is updated iteratively in Line 20 of Alg. 1, ensuring that the set of high-resistance areas remains relevant and accurately reflects the current sampling challenges. This iterative update is crucial for maintaining the effectiveness of the up-sampling strategy. Overall, the up-sampling of high-resistance areas is critical for the parameter tuner’s training as it enhances the planner’s robustness, promotes better generalization to new environments, and addresses training inefficiency.
@@ V Evaluation
Table I: Navigation performance of EffiTune and baseline methods under Same-Env setup.
Metric $\Rightarrow$ Tuner $\Downarrow$ Easy Medium Difficult
@@
Table II: Navigation performance of EffiTune and baselines under Cross-Env and Cross-Level setups. The column names represent the test environment levels: All (across all three levels), Easy, Medium and Difficult.
Training Env. Metric $\Rightarrow$ Tuner $\Downarrow$ All Easy Medium Difficult
@@
In evaluation, we answer research questions (RQ) below:
RQ1: How is EffiTune’s tuning performance compared with baseline tuning methods (under the same time budget)?
RQ2: Can EffiTune improve parameter tuner’s robustness under different environments and difficulty levels?
RQ3: How effectiveness is EffiTune’s diagnosis (region identification and failure filtering) for parameter tuners?
RQ4: Can EffiTune improve the efficiency of training process of parameter tuners?
RQ5: Can the advantage of EffiTune be illustrated through a case study?
@@ V-A Experiment Setup
Simulation & Planners. Following existing works [xiao2020appld, xu2021applr], we implement EffiTune using the widely-used Gazebo [Gazebo] simulator with a ClearPath Jackal differential drive ground robot as a robot model. Due to space limit, we leave the details of the classical planner and the parameter tuner that EffiTune builds upon for evaluation to Appendix I.
Baselines & Dataset. We compare EffiTune with the following tuning setups: (1) classical planners (DWA [fox1997dwa] without parameter tuner; (2) manually-tuned by expert (DWA-Fast [dwafast]); (3) one representative static parameter tuner, BayesOpt [snoek2012practical] and (4) one SOTA dynamic parameter tuner APPLR [xu2021applr]. We choose the same dataset as APPLR, namely BARN [perille2020benchmarking] dataset, as it includes various maps for robot navigation inside the Gazebo simulator and it has different difficulty levels’ map (i.e., Easy, Medium, and Difficult) to evaluate. This allows for a convenient evaluation of various planners’ performance across these varied settings.
@@
Comparison Setup. We introduce 3 evaluation setups: (1) Same-Env: The parameter tuner is trained and tested in the same environment set with predefined but different start positions to avoid overfitting. (2) Cross-Env: The parameter tuner is trained and tested at the same difficulty level but in different environments. (3) Cross-Level: The parameter tuner is trained on one difficulty level but tested on a different difficulty level.
V-B Same-environment Performance (RQ1)
We evaluate EffiTune and four baselines under the Same-Env setup. DWA, DWA-Fast, and BayesOpt are static planners, so we directly implement their fixed configurations from the training environment and test them on the same environments with random starting positions. For APPLR and EffiTune, we train the parameter tuner on the training set and deploy it on the test environments.
Table I shows that EffiTune consistently achieves the highest NS score across all difficulty levels, demonstrating its ability to identify and improve performance in underrepresented scenarios. While DWA performs better on SR and CR in easy and medium environments, EffiTune excels in the comprehensive NS metric. DWA’s conservative design prioritizes safety over time efficiency, leading to much longer navigation times, as seen in reduced ATT and lower NS scores, making it less practical and ideal for navigation tasks. Since NS is a widely-used metric [xu2021applr, xiao2020appld], EffiTune’s higher NS score effectively demonstrates its superiority over both static planners and the SOTA dynamic parameter tuner.
(a) Trained on Easy Level
(b) Trained on Medium Level
(c) Trained on Difficult Level
Figure 5: Comparison of RandomSampling and EffiTune. Parameter tuners are tested on all 3 levels.
V-C Cross-environment & Cross-level Performance (RQ2)
In this section, we compare EffiTune and selected baselines under the Cross-Env and Cross-Level setup, as shown in the diagonal sub-tables in Table II. Under the Cross-Env setup, EffiTune consistently achieves the highest NS score across all three environments, indicating its effectiveness in identifying and addressing underrepresented scenarios to enhance planner performance. For the Cross-Level experiments, EffiTune excels in generalization by achieving the highest NS scores when the trained parameter tuner is tested across different environmental difficulty levels. This demonstrates EffiTune’s robustness and capability to maintain navigation efficacy in new and varying conditions. Additionally, when parameter tuners trained in difficult environments are applied to easier ones, EffiTune outperforms even those baselines specifically trained for simpler environments, while this trend is not observed in the baseline methods. This result suggests that the parameter tuners trained by our method can learn more sophisticated and adaptable configurations, enabling EffiTune to excel in environments that are less challenging.
V-D Diagnosis Effectiveness (RQ3)
To better understand why our diagnosis strategy is effective, we introduce a variant called RandomSampling (RS). Unlike EffiTune, which identifies high-resistance areas with Alg. 2, RS indiscriminately samples from all locations reached by the robot during training without strategic guidance, while all other designs remain consistent with EffiTune. Results in Fig. 5 show that EffiTune significantly outperforms RS, with higher NS scores in all environments. The reasons are two-fold. First, RS fails to expose the robot to challenging, hard-to-reach scenarios, limiting effective training in these high-resistance areas, which is a crucial aspect for satisfactory navigation performance. Second, as shown in Fig. 2, medium and hard environments have more complicated maps with many hard-to-reach scenarios and obstacles than the easy level. Random sampling in such environments often leads to the robot encountering “dead-corner” scenarios – areas that are hard to exit, leading to extended, unrewarding training trajectories. Such trajectories will destabilize the training of parameter tuner and degrade navigation performance. This analysis verifies the necessity of our high-resistance area up-sampling step in EffiTune. By strategically selecting those hard-to-reach areas, we enhance the robot’s ability to perform well across a variety of navigational scenarios.
We also conduct an ablation study about failure trajectory filtering. Table III shows the impact of applying failure filtering on EffiTune’s performance metrics within the same computational budget. With filtering enabled, the Navigation Score (NS) improved to 23.23, demonstrating better navigation quality. Additionally, the Success Rate (SR) increased significantly from 70.59% to 76.47%, accompanied by a notable reduction in the Collision Rate (CR) from 22.06% to 17.65%. This indicates that the filtering process effectively removes noisy or less beneficial training data, enabling the parameter tuner to train more efficiently and enhancing overall robot navigation performance. Appendix II shows more results in the comparison between RS and EffiTune.
@@ V-D Diagnosis Effectiveness (RQ3)
V-E Training Efficiency (RQ4)
In this section, we evaluate the performance of different parameter tuner checkpoints during training under the Cross-Env setup when EffiTune is applied in a medium-level environment. Fig. 6 shows the average NS metric when those checkpoints are evaluated in testing environments. EffiTune consistently outperforms APPLR, achieving higher navigation scores at all training stages with the same training steps. Furthermore, APPLR requires over four times training steps to achieve comparable performance of EffiTune, and our method ultimately outperforms APPLR by 13.5%. These results demonstrate that our method significantly enhances the training efficiency of the parameter tuner and improves the overall navigation performance of the robot. Besides NS, we also show more metrics of different parameter tuner checkpoints during training (Appendix III).
Table III: Ablation Study. Performance of EffiTune with and without applying failure filtering.
@@ V-E Training Efficiency (RQ4)
V-F Case Study (RQ5)
In Fig. 7, we show a comparison before and after applying EffiTune to the parameter tuner. By leveraging our diagnosis and up-sampling techniques, the training process no longer always begins from the default position. Instead, it starts directly in the critical region, where the robot behavior exhibits signals of struggle and challenge. By skipping less beneficial training data and emphasizing the critical samples, we significantly boost the training efficiency of the parameter tuners.
(a) Original: Iter. 10 (Fig. 3(b))
(b) With EffiTune: Iter. 10
(c) Original: Iter. 90 (Fig. 3(d))
(d) With EffiTune: Iter. 90
Figure 7: Case Study for Illustration. The yellow arrow denotes the start position and the yellow circle denotes the goal.
V-G Ablation study
@@
We conduct ablation study to justify the importance of key components in EffiTune and the sensitivity test against the key hyper-parameters. Details can be found in Appendix IV.
@@ VI Conclusion
In this paper, we identify and analyze the training inefficiency of parameter tuners in robot navigation systems. To address these issues, we propose EffiTune, a diagnostic and mitigation framework for parameter tuners. EffiTune analyzes robot behaviors, identifies exploration bottlenecks and applies a targeted up-sampling strategy to balance the sampling data, enabling more effective learning and exploration. Our experiments demonstrate that EffiTune improves navigation performance by 13.5%+, enhances robustness in unseen environments, and quadruples training efficiency.
Limitations. Although EffiTune greatly boosts efficiency, further optimization of training duration is possible. The current pipeline depends on executing the classical planner to collect data, meaning increased complexity of the classical planner can raise training costs.
Acknowledgement
We are grateful to the Center for AI Safety for providing computational resources. This work was funded in part by the National Science Foundation (NSF) Awards SHF-1901242, SHF-1910300, Proto-OKN 2333736, IIS-2416835, DARPA VSPELLS - HR001120S0058, ONR N00014-23-1-2081, and Amazon. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsors.
@@ I Experiment Setup
The parameter tuner policy $\pi$ is trained to update the planning configurations $\theta$ of the classical planner. We maintain the same settings as APPLR, utilizing the DWA classical planner and the same action space $\theta$, including $\tt max\_vel\_x$}, $\tt max\_vel\_theta$}, $\tt vx\_samples$}, $\tt vtheta\_samples$}, $\tt occdist\_scale$}, $\tt pdist\_scale$}, $\tt gdist\_scale$}, and $\tt inflation\_radius$}. [^1] https://wiki.ros.org/dwa_local_planner We use the ROS dynamic reconfigure client to dynamically change planner parameters. The global goal $\alpha$ is assigned manually, while $\theta_{0}$ is the default set of parameters provided by the robot manufacturer. [^2] http://wiki.ros.org/costmap_2d Table IV shows the details of each configuration of the parameter tuner.
II Diagnosis Effectiveness
In this section, we delve deeper into the comparative analysis between RandomSampling and EffiTune by applying both methods across environments of varying complexity: Easy, Medium, and Difficult. We conducted separate experiments for each difficulty level, which allowed for detailed and direct comparisons. The results of these experiments are represented in Figures 8, 9, and 10.
Consistent with the observations discussed in Section V-D, our EffiTune continues to outperform RandomSampling in overall navigation performance. This superiority is quantified by a higher Navigation Score (NS), indicating EffiTune’s effectiveness in handling complex navigation tasks. This comparison not only highlights the strengths of EffiTune but also contributes to a better understanding of its operational dynamics across different environmental settings.